# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v1

@@ Abstract.
(1) we prove that the fixed loci of the natural anti-symplectic involutions on the moduli spaces of stable objects in the Kuznetsov component $\mathcal{K}u(X)$ of a Gushel–Mukai fourfold $X$ are 1-obstructed Lagrangian submanifolds, and
(2) we construct a family of immersed atomic Lagrangian submanifolds on each moduli space of stable objects in $\mathcal{K}u(X)$.
@@ 1. Introduction
Let $X$ be a projective hyper-Kähler manifold. The first effort in this direction is the theory of modular sheaves introduced by O’Grady in [55], whose discriminant satisfies a certain numerical equation. Then, it is shown in [55] that modular sheaves admit a wall-chamber decomposition in the ample cone and provide potential candidates for hyperholomorphic sheaves in the sense of [68]. Later, in [49], the author studies a class of objects of $\mathrm{D}^{\mathrm{b}}(X)$, called 1-obstructed objects (cf. Section 4.1.3), that can infinitesimally deform along a codimension one subspace of the second Hochschild cohomology $\mathrm{HH}^{2}(X)$. At the same time, for $F\in\mathrm{D}^{\mathrm{b}}(X)$, rather than $F$ itself, the author of [5] investigates when the Mukai vector of $F$ can infinitesimally deform along $\mathrm{HH}^{2}(X)$ in codimension one. If this is the case, such $F$ is called an atomic object (cf. Definition 4.1). By [49, Theorem 6.13], the extended Mukai vectors of atomic objects are as simple as the Mukai vectors of coherent sheaves on a K3 surface. Moreover, according to [5, Proposition 1.5], atomic torsion-free sheaves are modular.
@@
More precisely, for a projective hyper-Kähler manifold $X$, we say an immersed Lagrangian submanifold is $1$-obstructed or atomic if the push-forward of its structure sheaf is (cf. Definition 4.8). Assume that there exists a suitable connected atomic Lagrangian submanifold $L$ in $X$ with the first Betti number $b_{1}(L)>0$ and consider the component $\mathcal{M}$ of the moduli space of stable sheaves containing $\operatorname{\mathcal{O}}_{L}$. Then $\mathcal{M}$ contains a smooth open dense subscheme $\mathcal{M}^{0}$ that has a holomorphic symplectic structure. As discussed in [49, Section 7.4], a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}$ may yield a projective hyper-Kähler manifold with $b_{2}(\overline{\mathcal{M}})\geq b_{2}(X)+1$ and $\dim\overline{\mathcal{M}}=b_{1}(L)$.
@@ 1.1. Atomic Lagrangians in Bridgeland moduli spaces
  • When $n=3$ or $5$, there is a family of stability conditions on $\mathcal{K}u(X)$ constructed in [9], now called Serre-invariant stability conditions. The numerical K-group $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is a rank two lattice, generated by classes $\lambda_{1}$ and $\lambda_{2}$.
@@
In the situation (2)(a), as the hyperplane section $Y\subset X$ can vary in the linear series $|\operatorname{\mathcal{O}}_{X}(H)|$, we obtain an $8$-dimensional family of immersed atomic Lagrangians on each $M^{X}_{\sigma_{X}}(a,b)$ when $\dim X=4$, and a $10$-dimensional family when $\dim X=6$. On the other hand, in the situation (2)(b), we also get a $10$-dimensional family of immersed atomic Lagrangians on each $M^{X}_{\sigma_{X}}(a,b)$ when $Y$ varies in the moduli of GM fivefolds containing $X$ (cf. [34, Lemma 5.4]).
@@ 1.2. A non-rigid atomic Lagrangian submanifold
which is shown to be Lagrangian (cf. [34, Proposition 5.2, 5.6]). Moreover, these Lagrangians cover $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ when $X$ is fixed and $Y$ varies. See Section 6.1 for a review of these constructions.
It has been demonstrated in [35, 27] and Remark 6.12 that both of these varieties can be described in terms of Bridgeland moduli spaces. Furthermore, it is calculated in [19, Proposition 2.5] and [46, Proposition 0.5] that $\mathrm{H}^{1}(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2},\mathbb{C})=\mathbb{C}^{20}$. By applying Theorem 1.1, we establish:
@@
Thus, as explained above (see also Section 7.1), a hyper-Kähler compactification of an open dense subscheme $\mathcal{M}^{0}$ of the component of the moduli space of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing
@@
may yield a $20$-dimensional projective hyper-Kähler manifold of a new deformation type. Moreover, the support morphism of $\mathcal{M}^{0}$ to the Hilbert scheme of Lagrangians in $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is closely related to the relative Jacobian fibration of GM fivefolds containing $X$. We hope that the techniques and ideas in [48] can be applied to this case to construct a suitable compactification of $\mathcal{M}^{0}$. See Section 7.1 for a more detailed discussion.
@@ 1.3. Other examples
On the other hand, for a general cubic fourfold $X$, there are two families of hyper-Kähler manifolds associated with $X$: the Fano variety of lines $F(X)$ [4] and the LLSvS eightfold $\mathsf{Z}_{X}$ [45]. The hyper-Kähler fourfold $F(X)$ admits two different families of Lagrangian submanifolds. One of them is the Fano surface of lines in a general hyperplane section of $X$, which is atomic (cf. [49, Section 3], [5, Section 8]). Another family of Lagrangians is constructed from the Fano surface of planes in a general cubic fivefold containing $X$ (cf. [33]). In Proposition 7.4, we demonstrate that these Lagrangians are not atomic, but there exist atomic sheaves supported on them. For the LLSvS eightfold $\mathsf{Z}_{X}$, the Lagrangian submanifolds are constructed in [63]. Analogously, in Proposition 7.3, we prove that while these Lagrangians are not atomic, there exist atomic sheaves supported on them.
In [24, Conjecture A.1], the authors speculate a version of the construction of Lagrangians in the moduli spaces of stable objects in the Kuznetsov components of cubic fourfolds, analogous to the case of GM fourfolds. Recently, this conjecture has been proved for very general cubic fourfolds by [44, Theorem 8.2]. We expect that these Lagrangians are not atomic in general, but one can find atomic sheaves supported on them, as illustrated by the examples in Section 7.2.
@@
In Section 5, we first show that the fixed locus of the natural anti-symplectic involution on a moduli space of stable objects in the Kuznetsov component of a GM fourfold or sixfold is $1$-obstructed (cf. Theorem 5.5). Then, using this, we prove in Corollary 5.6 that the first Chern class of a moduli space of stable objects in the Kuznetsov component of a GM threefold or fivefold is proportional to the Bayer–Macrì divisor on it. This enables us to construct families of immersed atomic Lagrangians in each moduli space of stable objects in the Kuznetsov component of GM fourfolds or sixfolds (cf. Theorem 5.8 and Theorem 5.10).
@@
In Section 7, we first explain the relationship between the compactification of deformation spaces of atomic Lagrangians in double dual EPW sextics and the relative Jacobians of GM fivefolds (Section 7.1). Subsequently, in Section 7.2, we provide two examples of atomic sheaves supported on non-atomic Lagrangian submanifolds constructed from cubic hypersurfaces (cf. Proposition 7.3 and 7.4).
@@ Acknowledgements
It is our pleasure to thank Emanuele Macrì and Giovanni Mongardi for many insightful comments and suggesting Proposition 4.11, Proposition 7.3(2), and Proposition 7.4(2). We would also like to thank Alessio Bottini, Sasha Kuznetsov, Alex Perry, Laura Pertusi, and Xiaolei Zhao for many useful discussions. We are especially grateful to Qizheng Yin for many helpful conversations and numerous suggestions. HG would like to express sincere gratitude to Zhiyuan Li for his academic support and encouragement. ZL would like to thank the Institute for Advanced Study in Mathematics at Zhejiang University for financial support and wonderful research environment.
@@ 2.1. Kuznetsov components
By [15, Corollary 4.6], $X$ is non-Hodge-special when $X$ is very general among all ordinary GM varieties of dimension $n$ or very general among all special GM varieties of dimension $n$.
@@
When $n=3$, according to the proof of [41, Proposition 3.9], $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is a rank two lattice generated by $\lambda_{1}$ and $\lambda_{2}$, where
@@
where $\gamma^{*}_{X}\sigma_{2}$ is the pull-back of the Schubert cycle $\sigma_{2}\in\mathrm{H}^{4}(\operatorname{Gr}(2,5),\mathbb{Z})$. When $X$ is non-Hodge-special, by [38, Proposition 2.25], we have $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))=\langle\Lambda_{1},\Lambda_{2}\rangle$.
When $n=5$, by [39, Corollary 6.5], we can find a smooth GM threefold $X^{\prime}$ with an equivalence $\mathcal{K}u(X^{\prime})\simeq\mathcal{K}u(X)$. Hence, $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is also a rank two lattice with the Euler pairing (2).
When $n=6$, by [39, Corollary 6.5] again, we can find a smooth GM fourfold $X^{\prime}$ with an equivalence $\mathcal{K}u(X^{\prime})\simeq\mathcal{K}u(X)$. Hence, $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ contains a rank two lattice whose Euler pairing is the same as (4), which is the whole numerical Grothendieck group when $X$ is non-Hodge-special by [15, Corollary 4.6] and [38, Proposition 2.25].
We will use the following lemma. See e.g. [24, Lemma 4.11].
@@
We will also use the following relative version of Kuznetsov components. We say a smooth projective morphism $\pi\colon\mathcal{X}\to S$ is a family of GM varieties of dimension $n$ if there exists a relative ample line bundle $\operatorname{\mathcal{O}}_{\mathcal{X}}(1)$ such that $\mathcal{X}_{s}$ is a smooth GM variety of dimension $n$ and $\operatorname{\mathcal{O}}_{\mathcal{X}_{s}}(1)$ is the ample generator of $\operatorname{Pic}(\mathcal{X}_{s})$ for each $s\in S$ (cf. [20, Definition 3.1]). By [12, Lemma 5.9], when $S$ is noetherian and $n\geq 3$, there is a $S$-linear semiorthogonal decomposition
@@
By [8, Theorem 3.17, Lemma 3.25] or [56, Lemma 3.15], there is also a relative projection functor
@@ 2.2. Hodge structures of Kuznetsov components
By [2, Theorem 2.1] and [17, Proposition 3.4], the Mukai vector induces an embedding
@@
which is compatible with the pairing on both sides, up to a sign. Moreover, it is an isomorphism by [58, Proposition 8.2].
@@
When $n=4$, this follows from [57, Proposition 3.1]. When $n=6$, using (3) and the relations in Lemma 2.2, a similar argument as in [57, Proposition 3.1] yields the result. ∎
@@ 2.3. Stability conditions on Kuznetsov components
Indeed, in this case, we have $Z_{\sigma_{Y}}(j^{*}F)=Z_{\sigma_{X}}(\mathrm{pr}_{X}(j_{*}j^{*}F))$ by Lemma 2.2, then the claim follows from [24, Lemma 4.10(1), Lemma 4.3(2)].
@@
When $Y$ is a GM threefold, stability conditions on $\mathcal{K}u(Y)$ are constructed in [9]. It is proved in [61] that they are Serre-invariant (cf. [35, Definition 4.10]). Furthermore, they all belong to the same $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$-orbit [^1] The action of $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$ on stability conditions are defined in [13, Lemma 8.2]. Using the equivalence in [39, Corollary 6.5], the same result holds for GM fivefolds.
Theorem 2.7 ([35, Theorem A.10]).
@@
For non-Hodge-special GM varieties of even dimension, the following result is proved in [24, Proposition 4.12].

### v2

@@ Abstract.
(1) we prove that the fixed loci of the natural anti-symplectic involutions on the moduli spaces of stable objects in the Kuznetsov component $\mathcal{K}u(X)$ of a Gushel–Mukai fourfold $X$ are 1-obstructed Lagrangian submanifolds,
(2) we construct a family of immersed atomic Lagrangian submanifolds on each moduli space of stable objects in $\mathcal{K}u(X)$, and
[…]
@@ 1. Introduction
Let $X$ be a projective hyper-Kähler manifold. One effort in this direction is the theory of modular sheaves introduced by O’Grady in [61], whose discriminant satisfies a certain numerical equation. Then, it is shown in [61] that modular sheaves admit a wall-chamber decomposition in the ample cone and provide potential candidates for hyperholomorphic sheaves in the sense of [75]. Later, in [54], the author studies a class of objects of $\mathrm{D}^{\mathrm{b}}(X)$, called 1-obstructed objects (cf. Section 4.1.3), that can infinitesimally deform along a codimension one subspace of the second Hochschild cohomology $\mathrm{HH}^{2}(X)$. At the same time, for $F\in\mathrm{D}^{\mathrm{b}}(X)$, rather than $F$ itself, the author of [5] investigates when the Mukai vector of $F$ can infinitesimally deform along $\mathrm{HH}^{2}(X)$ in codimension one. If this is the case, such $F$ is called an atomic object (cf. Definition 4.1). By [54, Theorem 6.13], the extended Mukai vectors of atomic objects are as simple as the Mukai vectors of coherent sheaves on a K3 surface. Moreover, according to [5, Proposition 1.5], atomic torsion-free sheaves are modular.
@@
More precisely, for a projective hyper-Kähler manifold $X$, we say an immersed Lagrangian submanifold is $1$-obstructed or atomic if the push-forward of its structure sheaf is (cf. Definition 4.8). Assume that there exists a suitable connected atomic Lagrangian submanifold $L$ in $X$ with the first Betti number $b_{1}(L)>0$ and consider the component $\mathcal{M}$ of the moduli space of stable sheaves containing $\operatorname{\mathcal{O}}_{L}$. Then $\mathcal{M}$ contains a smooth open dense subscheme $\mathcal{M}^{0}$ that has a holomorphic symplectic structure. As discussed in [54, Section 7.4], a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}$ may yield a projective hyper-Kähler manifold with $b_{2}(\overline{\mathcal{M}})\geq b_{2}(X)+1$ and $\dim\overline{\mathcal{M}}=b_{1}(L)$.
@@ 1.1. Atomic Lagrangians in Bridgeland moduli spaces
  • When $n=3$ or $5$, there is a family of stability conditions on $\mathcal{K}u(X)$ constructed in [9], now called Serre-invariant stability conditions. The lattice $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is a rank two lattice, generated by classes $\lambda_{1}$ and $\lambda_{2}$.
@@
In the situation (2)(a), as the hyperplane section $Y\subset X$ can vary in the linear series $|\operatorname{\mathcal{O}}_{X}(H)|$, we obtain an $8$-dimensional family of immersed atomic Lagrangians on each $M^{X}_{\sigma_{X}}(a,b)$ when $\dim X=4$, and a $10$-dimensional family when $\dim X=6$. On the other hand, in the situation (2)(b), we also get a $10$-dimensional family of immersed atomic Lagrangians on each $M^{X}_{\sigma_{X}}(a,b)$ when $Y$ varies in the moduli of GM fivefolds containing $X$ (cf. [36, Lemma 5.4]).
[…]
Subsequently, based on [11, 12], starting from the atomic Lagrangians in Theorem 1.1, we obtain projectively hyperholomorphic twisted bundles on a series of hyper-Kähler manifolds.
Theorem 1.2.
@@ 1.2. A non-rigid atomic Lagrangian submanifold
which is shown to be Lagrangian (cf. [36, Proposition 5.2, 5.6]). Moreover, these Lagrangians cover $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ when $X$ is fixed and $Y$ varies. See Section 6.1 for a review of these constructions.
It has been demonstrated in [37, 28] and Remark 6.12 that both of these varieties can be described in terms of Bridgeland moduli spaces. Furthermore, it is calculated in [20, Proposition 2.5] and [49, Proposition 0.5] that $\mathrm{H}^{1}(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2},\mathbb{C})=\mathbb{C}^{20}$. By applying Theorem 1.1, we establish:
@@
As a corollary, the same argument as in Theorem 1.2 implies the following.
Corollary 1.4.
[…]
As explained above (see also Section 7.1), a hyper-Kähler compactification of an open dense subscheme $\mathcal{M}^{0}$ of the component of the moduli space of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing
@@
may yield a $20$-dimensional projective hyper-Kähler manifold of a new deformation type. Moreover, the support morphism of $\mathcal{M}^{0}$ to the Hilbert scheme of Lagrangians in $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is closely related to the relative Jacobian fibration of GM fivefolds containing $X$. We hope that the techniques and ideas in [51, 11, 12] can be applied to this case to construct a suitable compactification of $\mathcal{M}^{0}$. See Section 7.1 for a more detailed discussion.
@@ 1.3. Other examples
On the other hand, for a general cubic fourfold $X$, there are two families of hyper-Kähler manifolds associated with $X$: the Fano variety of lines $F(X)$ [4] and the LLSvS eightfold $\mathsf{Z}_{X}$ [48]. The hyper-Kähler fourfold $F(X)$ admits two different families of Lagrangian submanifolds. One of them is the Fano surface of lines in a general hyperplane section of $X$, which is atomic (cf. [54, Section 3], [5, Section 8]). Another family of Lagrangians is constructed from the Fano surface of planes in a general cubic fivefold containing $X$ (cf. [35]). In Proposition 7.5, we demonstrate that these Lagrangians are not atomic, but there exist atomic sheaves supported on them. For the LLSvS eightfold $\mathsf{Z}_{X}$, the Lagrangian submanifolds are constructed in [70]. Analogously, in Proposition 7.3, we prove that while these Lagrangians are not atomic, there exist atomic sheaves supported on them. As in Theorem 1.2 and Corollary 1.4, we also use these examples to construct stable atomic twisted bundles (cf. Corollary 7.4 and 7.6).
In [25, Conjecture A.1], the authors speculate a version of the construction of Lagrangians in the moduli spaces of stable objects in the Kuznetsov components of cubic fourfolds, analogous to the case of GM fourfolds. Recently, this conjecture has been proved for very general cubic fourfolds by [47, Theorem 8.2]. We expect that these Lagrangians are not atomic in general, but one can find atomic sheaves supported on them, as illustrated by the examples in Section 7.2.
@@
In Section 5, we first show that the fixed locus of the natural anti-symplectic involution on a moduli space of stable objects in the Kuznetsov component of a GM fourfold or sixfold is $1$-obstructed (cf. Theorem 5.5). Then, using this, we prove in Corollary 5.6 that the first Chern class of a moduli space of stable objects in the Kuznetsov component of a GM threefold or fivefold is proportional to the Bayer–Macrì divisor on it. This enables us to construct families of immersed atomic Lagrangians in each moduli space of stable objects in the Kuznetsov component of GM fourfolds or sixfolds (cf. Theorem 5.8 and Theorem 5.10). Finally, we prove Theorem 1.2 using these atomic Lagrangians.
@@
In Section 7, we first explain the relationship between the compactification of deformation spaces of atomic Lagrangians in double dual EPW sextics and the relative Jacobians of GM fivefolds (Section 7.1). Subsequently, in Section 7.2, we provide two examples of atomic sheaves supported on non-atomic Lagrangian submanifolds constructed from cubic hypersurfaces (cf. Proposition 7.3 and 7.5).
@@ Acknowledgements
It is our pleasure to thank Emanuele Macrì and Giovanni Mongardi for many insightful comments and suggesting Proposition 4.11, Proposition 7.3(2), and Proposition 7.5(2). We would also like to thank Alessio Bottini, Sasha Kuznetsov, Chunyi Li, Alex Perry, Laura Pertusi, and Xiaolei Zhao for many useful discussions. We are especially grateful to Qizheng Yin for many helpful conversations and numerous suggestions. HG would like to express sincere gratitude to Zhiyuan Li for his academic support and encouragement. ZL would like to thank the Institute for Advanced Study in Mathematics at Zhejiang University for financial support and wonderful research environment.
@@ 2.1. Kuznetsov components
By [16, Corollary 4.6], $X$ is non-Hodge-special when $X$ is very general among all ordinary GM varieties of dimension $n$ or very general among all special GM varieties of dimension $n$.
@@
When $n=3$, according to the proof of [44, Proposition 3.9], $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is a rank two lattice generated by $\lambda_{1}$ and $\lambda_{2}$, where
@@
where $\gamma^{*}_{X}\sigma_{2}$ is the pull-back of the Schubert cycle $\sigma_{2}\in\mathrm{H}^{4}(\operatorname{Gr}(2,5),\mathbb{Z})$. When $X$ is non-Hodge-special, by [41, Proposition 2.25], we have $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))=\langle\Lambda_{1},\Lambda_{2}\rangle$.
When $n=5$, by [42, Corollary 6.5], we can find a smooth GM threefold $X^{\prime}$ with an equivalence $\mathcal{K}u(X^{\prime})\simeq\mathcal{K}u(X)$. Hence, $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ is also a rank two lattice with the Euler pairing (2).
When $n=6$, by [42, Corollary 6.5] again, we can find a smooth GM fourfold $X^{\prime}$ with an equivalence $\mathcal{K}u(X^{\prime})\simeq\mathcal{K}u(X)$. Hence, $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(X))$ contains a rank two lattice whose Euler pairing is the same as (4), which is the whole numerical Grothendieck group when $X$ is non-Hodge-special by [16, Corollary 4.6] and [41, Proposition 2.25].
We will use the following lemma. See e.g. [25, Lemma 4.11].
@@
We will also use the following relative version of Kuznetsov components. We say a smooth projective morphism $\pi\colon\mathcal{X}\to S$ is a family of GM varieties of dimension $n$ if there exists a relative ample line bundle $\operatorname{\mathcal{O}}_{\mathcal{X}}(1)$ such that $\mathcal{X}_{s}$ is a smooth GM variety of dimension $n$ and $\operatorname{\mathcal{O}}_{\mathcal{X}_{s}}(1)$ is the ample generator of $\operatorname{Pic}(\mathcal{X}_{s})$ for each $s\in S$ (cf. [21, Definition 3.1]). By [13, Lemma 5.9], when $S$ is noetherian and $n\geq 3$, there is a $S$-linear semiorthogonal decomposition
@@
By [8, Theorem 3.17, Lemma 3.25] or [63, Lemma 3.15], there is also a relative projection functor
@@ 2.2. Hodge structures of Kuznetsov components
By [2, Theorem 2.1] and [18, Proposition 3.4], the Mukai vector induces an embedding
@@
which is compatible with the pairing on both sides, up to a sign. Moreover, it is an isomorphism by [65, Proposition 8.2].
@@
When $n=4$, this follows from [64, Proposition 3.1]. When $n=6$, using (3) and the relations in Lemma 2.2, a similar argument as in [64, Proposition 3.1] yields the result. ∎
@@ 2.3. Stability conditions on Kuznetsov components
Indeed, in this case, we have $Z_{\sigma_{Y}}(j^{*}F)=Z_{\sigma_{X}}(\mathrm{pr}_{X}(j_{*}j^{*}F))$ by Lemma 2.2, then the claim follows from [25, Lemma 4.10(1), Lemma 4.3(2)].
@@
When $Y$ is a GM threefold, stability conditions on $\mathcal{K}u(Y)$ are constructed in [9]. It is proved in [68] that they are Serre-invariant (cf. [37, Definition 4.10]). Furthermore, they all belong to the same $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$-orbit [^1] The action of $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$ on stability conditions are defined in [14, Lemma 8.2]. Using the equivalence in [42, Corollary 6.5], the same result holds for GM fivefolds.
Theorem 2.7 ([37, Theorem A.10]).
@@
For non-Hodge-special GM varieties of even dimension, the following result is proved in [25, Proposition 4.12].